# A new expansion of the Riemann zeta function

### v2

@@ 2.5 Interpolation of partial sums
Let $f\in\mathcal{E}^{\pi}$. There is a unique function $\varphi_{f}\in\mathcal{E}^{\pi}$, such that
@@ 3.6 An orthonormal expansion
$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$ (41)
@@
$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt$
$=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\frac{1}{2}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
@@
$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{n})=\log(2\pi)-\gamma-\frac{3}{2}.$
@@
With (42) and (44) this gives
$\int_{0}^{+\infty}\Big(\frac{1}{e^{t}-1}-\frac{1}{t}\Big)^{2}\ dt=\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2},$ (46)

### v3

@@ 2.5 Interpolation of partial sums
Let $f\in\mathcal{E}$. There is a unique function $\varphi_{f}\in\mathcal{E}$, such that
@@ 3.6 An orthonormal expansion
$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$
@@
$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \Gamma((\frac{1}{2}+it))\zeta(\frac{1}{2}+it)\ dt$
$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Corollary 2.
We have
$\gamma=1+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)},$
and for $m\geq 1$
$\zeta(m+1)=1+\frac{1}{m}+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{(\frac{1}{2}-it)_{m}}{m!}\frac{dt}{\cosh(\pi t)}.$
Proof.
With (41) and inversion of the binomial transform, we get
$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x_{k}=2^{m}X(m),$
and from (21)
$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{\frac{1}{2}P_{k}(\frac{1}{2}+it)}=2^{m}\ \frac{(\frac{1}{2}-it)_{m}}{m!}.$
Thus by (43) we obtain the integral expression
$X(m)$ $=$ $\frac{1}{2^{m+1}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\int_{-\infty}^{+\infty}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$=$ $\frac{1}{2^{m+1}}\int_{-\infty}^{+\infty}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\frac{(\frac{1}{2}-it)_{m}}{m!}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Example 8.
For $m=1,2,$ we have
$\zeta(2)$ $=$ $2+\frac{1}{2}\int_{-\infty}^{+\infty}(\frac{1}{2}-it)\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$\zeta(3)$ $=$ $\frac{3}{2}-\frac{1}{4}\int_{-\infty}^{+\infty}(t^{2}+2it-\frac{3}{4})\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
@@
$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{m})=\log(2\pi)-\gamma-\frac{3}{2}.$
@@
With (44) this gives
$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$ (46)